OpenStudy (anonymous):

MEDAL AWARD A cube has an edge which is 4 inches long. If the edge is increased by 25% then the volume is increased by approximately: (a) 25% (b) 48% (c) 73% (d) 97% (E)122%

OpenStudy (anonymous):

Let: V1 = original volume V2 = new volume S1 = original side length S2 = new side length S1 = 4 in V1 = length x width x height = 4 x 4 x 4 = 64 cubic inches S2 = [4 + (.25)4] in = (4 + 1) in = 5 in V2 = 5 x 5 x 5 = 125 cubic inches 64 + 64(97%) = 126.08 cubic inches The volume increased by approximately 97%.

OpenStudy (anonymous):

Oh. Thank you!